How do you figure out
x + 2/15 = 2/9 ????

Answer by checkley77(12844)

x + 2/15 = 2/9
X=2/9-2/15
X=(2*5-2*3)/45
X=(10-6)/45
X=4/45 ANS.
PROOF:
4/45+2/15=2/9
(4+2*3)/45=2/9
(4+6)/45=2/9
10/45=2/9
90=90